# EVCap: Retrieval-Augmented Image Captioning with External Visual–Name Memory for Open-World Comprehension

### 27 Nov 2023 v1

@@ 6 Conclusion
We further advance image captioning in real-world scenarios by introducing EVCap, a novel image captioning model with object names retrieved from an external visual–name memory. The external memory is easily expandable, allowing for effortless updates with new object visuals and names. EVCap stands out for its efficiency, comprising merely 3.97M trainable parameters, yet delivering robust performance. We extensively compare EVCap with SOTAs on various benchmarks and commonsense-violating data, demonstrating its significant superiority in performance.
@@
Figure A: Samples of images and corresponding names in the external visual–name memory constructed from LVIS’s objects.
@@ A Implementation Details
Our method is based on Pytorch and is trained within one epoch with a batch size of 24 using mixed precisions. We optimize the model using AdamW, setting the weight decay at 0.05, and using $\beta_{1}$ and $\beta_{2}$ values of 0.9 and 0.99, respectively. A cosine learning rate (LR) decay strategy is adopted, starting with an initial LR of 1e-4. The model undergoes 5000 linear warm-up steps, beginning with a start LR of 1e-6. During the evaluation phase, we use a beam search strategy with a beam size of 5 to generate captions.
Figure B: Samples of images and corresponding names in the added WHOOPS’s objects.
@@ B.1 LVIS memory
As stated in Sec. 3.2 of the main paper, we utilize 1203 objects from the LVIS dataset. For each of these objects, we randomly select between one and ten images from LVIS. Additionally, we enrich our data by incorporating five synthetic images for each object, created using stable diffusion. We show two samples of this external visual-name memory, constructed using objects from LVIS in Fig. A.
@@ B.2 WHOOPS memory
To illustrate the scalability of the external memory in EVCap, we expand it by integrating WHOOPS knowledge into the original external visual–name memory in Sec. 5.2 and Sec. 5.3 of the main paper. Specifically, we focus on objects that are mentioned in the answers of VQA annotations in the WHOOPS dataset because of their conciseness and emphasis on key objects. For each of these objects, we produce five synthetic images employing stable diffusion. Two examples from this augmented memory, featuring newly added object images and their corresponding names, are presented in Fig. B.
@@ C.1 Experiments on NoCaps test set
Qualitative results. Fig. D shows the captions generated by our EVCap alongside those from three SOTA methods on the NoCaps test set. It also includes the object names retrieved by EVCap and the captions retrieved by SmallCap. Consistent with the findings in the main paper, SmallCap tends to generate hallucinatory objects that are absent in the input images, such as “tie” and “mouse”. The same hallucinatory object “mouse” is also found in the retrieved captions, indicating that SmallCap’s diminished performance is largely due to its reliance on retrieved captions containing irrelevant information. In comparison, our EVCap demonstrates a performance on par with BLIP-2.
Figure C: CIDEr scores after training EVCap with the number of object name query tokens $P$ from 2 to 10. The results indicate that the performance is relatively optimal when $P$ is set to be 8.
Figure D: Examples of captions generated by our EVCap, and three SOTAs on the NoCaps test set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
@@ C.2 Further analysis
Number of object name query tokens. We explore the impact of varying the number of retrieved object names $P$ (Sec. 3.4 of the main paper) on EVCap in Fig. C. We train the model using different values of $P$, ranging from 2 to 10, and evaluate the performance under CIDEr on all three benchmarks. The results suggest that setting $P=8$ offers relatively optimal results.
Figure E: Examples of captions generated by our EVCap, and three SOTAs on the COCO test set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
Figure F: Examples of captions generated by our EVCap, and three SOTAs on the NoCaps validation set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
@@
More qualitative examples on the COCO test set, NoCaps validation set, and Flickr30k test set are shown in Fig. E, Fig. F, and Fig. G, respectively.
Figure G: Examples of captions generated by our EVCap, and three SOTAs on the Flickr30k test set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.

### 07 Apr 2024 v2

@@ 6 Conclusion
We further advance image captioning in real-world scenarios by introducing EVCap, a novel image captioning model with object names retrieved from an external visual–name memory. The external memory is easily expandable, allowing for effortless updates with new object visuals and names. We extensively compare EVCap with SOTAs on various benchmarks and commonsense-violating data, demonstrating its significant superiority in performance.
Acknowledgements. This work was supported by JSPS/MEXT KAKENHI Grant Numbers JP22H05015, JP23H03449, JP23KJ0404, and JP22K17947, and the commissioned research (No. 225) by the National Institute of Information and Communications Technology (NICT), Japan, and the Institute of AI and Beyond of the University of Tokyo, and ROIS NII Open Collaborative Research 2023-23FC01, 2024-24S1201.
@@ A Implementation Details
Customized Q-Former. Fig. A depicts customized Q-Former, where the image embedding port (CLIP visual feature port) now receives retrieved object names $\mathcal{S}$, while the text port receives visual features $\mathcal{Q}$.
Figure A: Customized Q-Former and original Q-Former [25].
Implementation. Our method is based on Pytorch and is trained within one epoch with a batch size of 24 using mixed precisions. We optimize the model using AdamW, setting the weight decay at 0.05, and using $\beta_{1}$ and $\beta_{2}$ values of 0.9 and 0.99, respectively. A cosine learning rate (LR) decay strategy is adopted, starting with an initial LR of 1e-4. The model undergoes 5000 linear warm-up steps, beginning with a start LR of 1e-6. During the evaluation phase, we use a beam search strategy with a beam size of 5 to generate captions.
Figure B: Samples of images and corresponding names in the external visual–name memory constructed from LVIS’s objects.
Figure C: Samples of images and corresponding names in the added WHOOPS’s objects.
@@ B.1 LVIS memory
As stated in Sec. 3.2 of the main paper, we utilize 1203 objects from the LVIS dataset. For each of these objects, we randomly select between one and ten images from LVIS. Additionally, we enrich our data by incorporating five synthetic images for each object, created using stable diffusion. We show two samples of this external visual-name memory, constructed using objects from LVIS in Fig. B.
@@ B.2 WHOOPS memory
To illustrate the scalability of the external memory in EVCap, we expand it by integrating WHOOPS knowledge into the original external visual–name memory in Sec. 5.2 and Sec. 5.3 of the main paper. Specifically, we focus on objects that are mentioned in the answers of VQA annotations in the WHOOPS dataset because of their conciseness and emphasis on key objects. For each of these objects, we produce five synthetic images employing stable diffusion. Two examples from this augmented memory, featuring newly added object images and their corresponding names, are presented in Fig. C.
@@ C.1 Experiments on NoCaps test set
Qualitative results. Fig. E shows the captions generated by our EVCap alongside those from three SOTA methods on the NoCaps test set. It also includes the object names retrieved by EVCap and the captions retrieved by SmallCap. Consistent with the findings in the main paper, SmallCap tends to generate hallucinatory objects that are absent in the input images, such as “tie" and “mouse". The same hallucinatory object “mouse" is also found in the retrieved captions, indicating that SmallCap’s diminished performance is largely due to its reliance on retrieved captions containing irrelevant information. In comparison, our EVCap demonstrates a performance on par with BLIP-2.
Figure D: CIDEr scores after training EVCap with the number of object name query tokens $P$ from 2 to 10. The results indicate that the performance is relatively optimal when $P$ is set to be 8.
Figure E: Examples of captions generated by our EVCap, and three SOTAs on the NoCaps test set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
@@ C.2 Further analysis
Number of object name query tokens. We explore the impact of varying the number of retrieved object names $P$ (Sec. 3.4 of the main paper) on EVCap in Fig. D. We train the model using different values of $P$, ranging from 2 to 10, and evaluate the performance under CIDEr on all three benchmarks. The results suggest that setting $P=8$ offers relatively optimal results.
Figure F: Examples of captions generated by our EVCap, and three SOTAs on the COCO test set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
Figure G: Examples of captions generated by our EVCap, and three SOTAs on the NoCaps validation set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.
@@
More qualitative examples on the COCO test set, NoCaps validation set, and Flickr30k test set are shown in Fig. F, Fig. G, and Fig. H, respectively.
Figure H: Examples of captions generated by our EVCap, and three SOTAs on the Flickr30k test set. We also list the retrieved object names by our EVCap (below the captions of EVCap) and retrieved captions by SmallCap (right side) in italics.